OpenStudy (anonymous):

what is the equation of the line symmetry for f(x)=2x2-12x-20?

OpenStudy (anonymous):

for \[y=ax^2+bx+c\]use \[x=-\frac{b}{2a}\]

OpenStudy (jchick):

Simplify 2x^2 - 12x

OpenStudy (jchick):

Equation becomes (2x^2 - 12x) - 20 = 0

OpenStudy (jchick):

Simplify 2x^2-12x - 20

OpenStudy (anonymous):

So, simplify the equation

OpenStudy (jchick):

Pull out like factors :

OpenStudy (jchick):

2x^2 - 12x - 20 = 2 • (x^2 - 6x - 10)

OpenStudy (jchick):

Factoring x^2 - 6x - 10

OpenStudy (jchick):

The first term is, x2 its coefficient is 1 . The middle term is, -6x its coefficient is -6 . The last term, "the constant", is -10

OpenStudy (jchick):

Multiply the coefficient of the first term by the constant 1 • -10 = -10

OpenStudy (jchick):

Find two factors of -10 whose sum equals the coefficient of the middle term, which is -6 .

OpenStudy (anonymous):

oh okay

OpenStudy (jchick):

-10 + 1 = -9 -5 + 2 = -3 -2 + 5 = 3 -1 + 10 = 9

OpenStudy (jchick):

2 • (x^2 - 6x - 10) = 0

OpenStudy (jchick):

Solve 2•(x^2-6x-10) = 0

OpenStudy (jchick):

Solve : 2 = 0

OpenStudy (jchick):

Is that true or false?

OpenStudy (anonymous):

I know that 2 does not equal 0

OpenStudy (jchick):

Ok so is it true or false?

OpenStudy (anonymous):

false

OpenStudy (jchick):

If true than it has a solution if not than no solutions.

OpenStudy (jchick):

This equation has no solution. A a non-zero constant never equals zero.

OpenStudy (anonymous):

yep, it does not

OpenStudy (jchick):

Do you need the vertex as well?

OpenStudy (anonymous):

well, sure if you want to

OpenStudy (anonymous):

Not sure what you're doing here, but this question doesn't ask for solutions to the quadratic. It asks for the line of symmetry, in other words the vertical line that passes through the vertex. @satellite73 showed you how to do it in the first reply to this post.

OpenStudy (anonymous):

oh yea, that's what i was about to say

OpenStudy (jchick):

Ok so from here we will find the vertex

OpenStudy (jchick):

The vertex of the parabola relates to its axis of symmetry and how to determine the axis of symmetry from a quadratic equation.

OpenStudy (jchick):

Find the Vertex of y = x^2-6x-10

OpenStudy (anonymous):

i'm trying to get question done, cuz i'm on my quiz, and i don't it want it to take long

OpenStudy (jchick):

Parabolas have a highest or a lowest point called the Vertex .

OpenStudy (jchick):

Ok well I was showing you how I do it but my way is different from others. We will end up at the same end point.

OpenStudy (anonymous):

You don't need solutions, you don't need the coordinates of the vertex. To answer the question, compare the given quadratic to \[y=ax^2 + bx + c\]Determine the values os a, b, and c. The equation of the line of symmetry is then\[x=-\frac{ b }{ 2a }\]

OpenStudy (anonymous):

okay so 2x2 is a, -12 is b and -20 is c?

OpenStudy (anonymous):

-12x i met for b

OpenStudy (anonymous):

a=2, b=-12, c=-20

OpenStudy (jchick):

@ospreytriple we will end up with the same answer I just do it differently.

OpenStudy (anonymous):

okay, so i think my answer is -3

OpenStudy (anonymous):

There's an issue with the sign. Check your work again very carefully.

OpenStudy (anonymous):

\[x=-\frac{ b }{ 2a } = -\frac{ -12 }{ 2(2) }\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

oh wait it's +3 is it?

OpenStudy (anonymous):

Right. But you have yo write the full equation of the line, i.e. \(x=3\)

OpenStudy (anonymous):

oh x=3 got it

OpenStudy (anonymous):

Good job.

OpenStudy (anonymous):

well, thank you very much, and other two help me good too